OpenStudy (anonymous):

How to simplify -cos^2xcos(pi/2-x)

OpenStudy (anonymous):

\[-\cos ^2x \cos \left( \frac{ \pi }{ 2 }-x \right)=-\cos^2 x \sin x=-\left( 1-\sin ^2x \right)\sin x=\sin x \left( \sin ^2x-1 \right)\]

OpenStudy (gorv):

-cos^2x*cos(pi/2-x)=-cosx*cosx*cos(pi/2-x) and cos(pi/2-x)=sinx

OpenStudy (gorv):

2cosx*sinx=sin2x cosx*sinx=(sin2x)/2

OpenStudy (gorv):

-(cosx*sin2x)/2 use sin C & D formula now